OpenStudy (anonymous):

What is the slope of the graph of y=-2/5x+3? And I also need help with the problem below. The length of a painting is two less than three times the width. If the width of the painting is 15 inches, what is the length in inches?

OpenStudy (anonymous):

well, y=mx+b, where m=slope. so in y=-2/5x+3, m=slope or -2/5=slope. the problem: l+2=2w w=15 l+2=2(15) l=28 therefore length=28in.